QUESTION 3 dy dx E-2xy2 and y=3 when x = 1. What is the value ofy when x = 3? 3 23 1 3 0 o w celu 28
Answer #1

dy - 2 152 3 1 / 키 2 ) 얘 dy =-2x da 화 dy -2facute C gig033 -놓544 키 오마층권 A 10 - 마 f 23 | 어느임 190 여 거23 = 9-38 2 25 3 2 27-2 3For any doubt please ask me..and if you satisfied please thumbs-up for this answer